# Boundary Time Crystals as AC sensors: enhancements and constraints

### v2

@@ 1 Introduction
We study the performance of the BTC by computing the quantum Fisher information (QFI) [27] of the system in response to changes in the strength of the applied AC field; a larger QFI implies an improved sensor performance. Here we consider solely the detection of weak signals; specifically, we work in the linear response limit where the strength of the signal to measured is vanishingly small. There are various bounds on the QFI depending on factors such as the class of parameter being sensed [28, 29], the structure of the Liouvillian [30, 31, 32], and correlations, or lack thereof, within the system [33]. These bounds are expressed in terms of the resources being utilised during the sensing protocol, in our case the relevant resources are the number of sites $N$ and the encoding time $t$. Given an open system sensing a time-dependent field, we rely on quite general bounds contingent on simply the presence of coherence and/or many-body correlations. For a classical system, with neither of these, the QFI is bounded by $Nt$; allowing for coherence increases this bound to $Nt^{2}$; and including many-body correlations gives us the Heisenberg limit: $N^{2}t^{2}$.
@@ 2 Model
where the $N$ spins are driven coherently with strength $\omega_{0}$ and decay collectively with a rate of $\kappa$ rescaled by $S=N/2$. We have introduced collective spin operators $\hat{S}^{\alpha=x,y,z}=\sum_{i}\hat{\sigma}^{\alpha}_{i}/2$ where $\hat{\sigma}^{\alpha}_{i}$ is the Pauli-$\alpha$ matrix acting on the $i$-th spin. These also describe a spin of length $S$ with raising/lowering operators $\hat{S}^{\pm}=\hat{S}^{x}\pm i\hat{S}^{y}$
In the thermodynamic ($N\to\infty$) limit, this model undergoes a dynamical phase transition as a function of $\omega_{0}/\kappa$: when $\omega_{0}/\kappa<1$ the spins decay to a steady state but for $\omega_{0}/\kappa>1$ they oscillate in perpetuity. This non-trivial phase corresponds to a breaking of the continuous time-translation symmetry of Eq. (2), occurring only in the thermodynamic limit and, in this way supporting a time crystal phase. Specifically this is known as a boundary time crystal (BTC), referring to the fact that the master equation can be derived as an effective description of $N$ spins on the boundary of a much larger closed system.
@@
Here, the AC field is characterized by its frequency, $\omega_{\text{ac}}$, and its phase shift, $\phi$. We have also introduced $\hat{\rho}_{g}(t)$: the density matrix conditioned on the value of the field strength, $g$, evolved to time $t$.
@@
To determine the sensitivity of the BTC to variations in the strength of this field $g$ we compute the quantum Fisher information (QFI). The QFI bounds the uncertainty in estimating the value of $g$; a larger QFI magnitude indicates an improved sensor performance. This is given by
$F_{g}(\hat{\rho})=\mathrm{Tr}[\hat{L}_{g}^{2}\hat{\rho}_{g}(t)],$ (4)
where we have introduced the symmetric logarithmic derivative (SLD), $L_{g}$, of a density matrix $\rho_{g}$, depending on a real parameter $g$, defined as
$\partial_{g^{\prime}}\hat{\rho}_{g^{\prime}}(t)\big{|}_{g^{\prime}=g}=\frac{1}{2}\{\hat{L}_{g},\hat{\rho}_{g}(t)\}.$ (5)
Expressing the SLD in the eigenbasis of $\hat{\rho}_{g}(t)$, with {|$\{\ket{\lambda_{i}}\}$ its eigenvectors, allows one to solve this equation. Substituting this solution into Eq. (4) yields [36, 37]
@@
where $\{\lambda_{i}\}$ are the eigenvalues of $\hat{\rho}_{g}$. Here we have suppressed the time dependence for conciseness. Computing $\partial_{g}\hat{\rho}_{g}$ can be done numerically by either a finite difference in terms of $g$ or, as we shall show, by computing a second master equation in conjunction with Eq. (2). We can find the time evolution of $\partial_{g}\hat{\rho}_{g}$ by simply taking the time derivative of Eq. (2), this given by
@@ 3 QFI Dynamics
In this section, we study the scaling of the QFI as a function of $N$ and $t$. We use its scaling behaviour to characterize the performance of the sensor and compare this performance with the limits of a classical sensor as well as the fundamental “Heisenberg” quantum limit. These limits are expressed in terms of the total resources used in the sensing protocol. Incorrect characterization of the resources used or omitting a certain resource, such as sensing time, can lead to results which seemingly surpass even the Heisenberg limit. We shall be careful to state which resources we consider in each result. All results presented here correspond to an initial state for which $\langle\hat{S}^{z}\rangle=S$, they were were generated using QuantumOptics.jl [38] and code adapted from this framework.
@@
The dynamics of the QFI (scaled by $N$) for various $N$ is plotted in Figure 2(a). These dynamics follow the typical behaviour of the QFI for a Markovian open quantum system [35, 39]: it initially increases over time before reaching some peak value and decaying. This initial growth arises from the coherence in the direction of the applied field accruing a phase. The coherence, however, only survives for a finite time due to the system being open and as this decays so does the ability to encode information. Consequently the QFI ceases growth, reaching a peak and subsequently decaying; we find that the time this peak occurs is related to the lifetime of the underlying time crystal dynamics.
@@ 4 Understanding the QFI Scaling
where we neglect terms of order smaller than O(S). The total density matrix in this limit can be described by the factorized ansatz as

### v3

@@ 1 Introduction
We study the performance of the BTC by computing the quantum Fisher information (QFI) [27] of the system in response to changes in the strength of the applied AC field; a larger QFI implies an improved sensor performance. Here we consider solely the detection of weak signals; specifically, we work in the linear response limit where the strength of the signal to measured is vanishingly small. There are various bounds on the QFI depending on factors such as the class of parameter being sensed [28, 29], the structure of the Liouvillian [30, 31, 32], and correlations, or lack thereof, within the system [33]. These bounds are expressed in terms of the resources being utilised during the sensing protocol, in our case the relevant resources are the number of spins $N$ and the encoding time $t$. Given an open system sensing a time-dependent field, we rely on quite general bounds contingent on simply the presence of coherence and/or many-body correlations. For a classical system, with neither of these, the QFI is bounded by $Nt$; allowing for coherence increases this bound to $Nt^{2}$; and including many-body correlations gives us the Heisenberg limit: $N^{2}t^{2}$.
@@ 2 Model
where the $N$ spins are driven coherently with strength $\omega_{0}$ and decay collectively with a rate of $\kappa$ rescaled by $N/2$. We have introduced collective spin operators $\hat{S}^{\alpha=x,y,z}=\sum_{i}\hat{\sigma}^{\alpha}_{i}/2$ where $\hat{\sigma}^{\alpha}_{i}$ is the Pauli-$\alpha$ matrix acting on the $i$-th spin. These also describe a spin of length $S$ with raising/lowering operators $\hat{S}^{\pm}=\hat{S}^{x}\pm i\hat{S}^{y}$
In the thermodynamic ($N\to\infty$) limit, this model undergoes a dynamical phase transition as a function of $\omega_{0}/\kappa$: when $\omega_{0}/\kappa<1$ the spins decay to a steady state but for $\omega_{0}/\kappa>1$ they oscillate in perpetuity. This non-trivial phase corresponds to a breaking of the continuous time-translation symmetry of Eq. (2), occurring only in the thermodynamic limit and, in this way supporting a time crystal phase. Specifically this is known as a Boundary Time Crystal (BTC), or alternatively as a dissipative time crystal, referring to the fact that the spins are only a subsystem of the total system-environment Hilbert space. Despite being a portion of the whole system, the boundary spins can be macroscopically large, and stabilise different phases of matter. Importantly, although the total system-environment Hamiltonian is time-translation invariant —consistent with various different derivations of the above effective master equation (see, e.g., supplementary information of Ref.[15] or Refs.[36, 37, 38]) — the boundary spins can still break this symmetry, leading to persistent oscillatory behaviour and thus supporting a time crystal phase.
@@
Here, the AC field is characterized by its frequency, $\omega_{\text{ac}}$, and its phase shift, $\phi$. We have also introduced $\hat{\rho}_{g}(t)$: the density matrix conditioned on the value of the field strength, $g$, evolved to time $t$. Our observations indicate that choosing the external field along the z-direction leads to higher values of sensor performance as compared to the x- or y-direction, making it a more effective choice for the sensing protocol. Although we have no formal analytical proof that this direction is always optimal, the key insight stems from the instability of the time crystal phase - or its higher susceptibility - to perturbations along the z-direction. This instability has been demonstrated in analytical studies within a mean-field treatment [39] and observed in generalised versions of the model [40, 41]. As a consequence, the system exhibits a stronger response to the applied AC field, thus allowing a greater precision in its estimation.
@@
To determine the sensitivity of the BTC to variations in the strength of this field $g$ we compute the quantum Fisher information (QFI). The QFI bounds the uncertainty in estimating the value of $g$, with $\Delta g\geq 1/\sqrt{MF(\hat{\rho}_{g}(t))}$, where $F(\hat{\rho}_{g}(t)$ is the QFI of the sensor probe, and $M$ is the number of measurements performed in the system in order to extract and therefore decode the information about $g$. Clearly, a larger QFI magnitude indicates an improved sensor performance. This is given by [42, 43]
@@
where we have introduced the density matrix $\rho_{g}$, depending on a real parameter $g$ and $\{\lambda_{i}\}$ are the eigenvalues of $\hat{\rho}_{g}$. Here we have suppressed the time dependence for simplicity.
Computing $\partial_{g}\hat{\rho}_{g}$ can be done numerically by either a finite difference in terms of $g$ or, as we shall show, by computing a second master equation in conjunction with Eq. (2). We can find the time evolution of $\partial_{g}\hat{\rho}_{g}$ by simply taking the time derivative of Eq. (2), this given by
@@
Some properties of the QFI are worth recalling. While for classical systems the QFI is limited simply by number of repetitions in the estimation protocol, leading to a linear scaling in time (one can think of partitioning the whole time into small time bins representing the number of measurements), exploiting quantum coherence in the system allows a quadratic improvement over time [44], extending the maximun of QFI to $F_{g}(\hat{\rho})<t^{2}$. Furthermore, since Fisher Information is additive, a sensor composed of $N$ separable spins obeys the bound $F_{g}(\hat{\rho})<Nt^{2}$. Surpassing this bound requires nonseparability, that is, harnessing quantum correlations among the spins within the sensor. The ultimate quantum advantage emerges when entanglement is fully exploited, allowing a quadratic gain in both the number of spins and time, thus achieving the Heisenberg limit with $F_{g}(\hat{\rho})=N^{2}t^{2}$.
@@ 3 QFI Dynamics
In this section, we study the scaling of the QFI as a function of $N$ and $t$. We use its scaling behaviour to characterize the performance of the sensor and compare this performance with the limits of a classical sensor as well as the fundamental “Heisenberg” quantum limit. These limits are expressed in terms of the total resources used in the sensing protocol. Incorrect characterization of the resources used or omitting a certain resource, such as sensing time, can lead to results which seemingly surpass even the Heisenberg limit. We shall be careful to state which resources we consider in each result. All results presented here correspond to an initial state for which $\langle\hat{S}^{z}\rangle=N/2$, they were generated using QuantumOptics.jl [45] and code adapted from this framework.
@@
The dynamics of the QFI (scaled by $N$) for various $N$ is plotted in Figure 2(a). These dynamics follow the typical behaviour of the QFI for a Markovian open quantum system [35, 46]: it initially increases over time before reaching some peak value and decaying. This initial growth arises from the coherence in the direction of the applied field accruing a phase. The coherence, however, only survives for a finite time due to the system being open; as it decays, so does the ability to encode information. Consequently the QFI ceases growth, reaching a peak and subsequently decaying; we find that the time this peak occurs is related to the lifetime of the underlying time crystal dynamics.
@@ 4 Understanding the QFI Scaling
where we neglect terms of order smaller than O(N). The total density matrix in this limit can be described by the factorized ansatz as
@@ Acknowledgements
Data availability:
The data and codes of this manuscript are available from the corresponding author upon reasonable request.